OpenStudy (anonymous):

solve for the variable 20^(-6n)=49

OpenStudy (anonymous):

20^(-6n)+6=55 I need help with going through the steps

OpenStudy (anonymous):

20^(-6n)=49 -6nlog(20)=log(49)

OpenStudy (anonymous):

THat is what I got so far

OpenStudy (anonymous):

Can you please help?

OpenStudy (anonymous):

@midhun.madhu1987 can you help?

OpenStudy (anonymous):

Can someone help with me understanding what I need to do? I know that the objective is to eliminate the exponents, but I also know some ways in which this becomes -6n=log(49)/log(20)

OpenStudy (anonymous):

Is there a way to simplify this more without using calculators?